40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 8 of 58  Not logged in ELOG logo
ID Date Author Type Category Subject
  2525   Tue Dec 22 10:42:17 2020 ranaSummaryGeneralInstrument loan

I was thinking about getting this new current pre-amp from NF:


It seems to have a good noise performance and has a built in low pass filter and also a remote interface.

The FEMTO seems less fancy, but their noise performance is actually 2-3x better.


FEMTO DLPCA200 low noise preamp (brand new)

Keithley Source Meter 2450 (brand new) => Returned 11/23/2020

were brought to the OMC lab for temporary use.


  2524   Wed Dec 16 22:08:36 2020 anchalUpdateEquipment transferTook delay line box, compressed nitrogen cylinder and lens to 2um (Crackle) lab
  • Took a delay line box DB64 from QIL from the WOPO table to the 2um (formerly known as Crackle) lab. The box was marked Crackle on it.
  • Took the compressed nitrogen cylinder for optics cleaning which was stored in Adaptive Optics lab.
  • Took some lens from the cabinet in Adaptive optics lab.
  • Took some other optics parts like pedestals, posts, lens mount etc.

See SUS_Lab/1877

  2523   Tue Dec 8 11:27:51 2020 RadhikaMiscLab MonitoringParticle counter communication

I have a python script for communication with the Met One 227a particle counter, but it appears like I am not receiving a response from the device. I swapped out a few serial-USB cables, but this did not fix the issue. I suspect I am making an unsound assumption about the communication protocol so I am going back to basics and reading more on RS232 and ASCII commands. 

  2522   Fri Nov 20 18:49:43 2020 KojiSummaryGeneralInstrument loan

FEMTO DLPCA200 low noise preamp (brand new)

Keithley Source Meter 2450 (brand new) => Returned 11/23/2020

were brought to the OMC lab for temporary use.

  2521   Fri Nov 20 18:47:42 2020 KojiSummaryGeneralPermenant exchange of TED200C(QIL) and TED200C(2umECDL)

I moved the brand new TED200C on the workbench to Crackle for 2um ECDL (permanently)
The TED200C temp controller used in the 2um PD test setup will stay there (permanently)


  2520   Thu Nov 19 16:25:30 2020 AidanUpdate2um PhotodiodesAgilis piezo mirror installed in JPL PD testing apparatus

Here's the python code I used to control this. 

I incorrectly used the Move to Limit command ('1MV-3': axis 1, MoVe, negative direction, speed 3', where the speeds are given in the manual, see Section 4.7 in particular). Once this command is issued, the stage will keep moving until it receives the stop command. The JOG command would be more appropriate.

I confirmed a smooth change in the PD output as the beam translated across it.


I installed the Agilis mirror before the lens and cryo-chamber. Used the USB interface to align the beam onto the PD. So we can control the alignment remotely now (or once I’ve properly connected the USB cable instead of today’s janky test connection).


Attachment 1: Screen_Shot_2020-11-19_at_4.27.17_PM.png
  2519   Tue Nov 17 17:51:44 2020 AidanUpdate2um PhotodiodesAgilis piezo mirror installed in JPL PD testing apparatus

I installed the Agilis mirror before the lens and cryo-chamber. Used the USB interface to align the beam onto the PD. So we can control the alignment remotely now (or once I’ve properly connected the USB cable instead of today’s janky test connection).

Attachment 1: IMG_9401.jpg
Attachment 2: IMG_9400.jpg
  2518   Wed Nov 11 16:09:22 2020 anchalSummaryECDLAUX wavelength finesee requirements in mariner - Added Excess Scatter Noise

An issue was raised with last calculation about the fact that our sensing of PDH signal isn't ideal and in the real world there is scattering, clipping extra adding excess noise in the PDH loop. This noise primarily comes by the intensity noise imparted on promptly reflected light from the cavity via various shaking optics etc on the table before it goes to the PDH reflection RF photodiode.

This noise's coupling to the PDH loop is identical to how shot noise of light couples into the PDH loop i.e.:

  • Intensity noise of light is converted into voltage oise by the PDH photo diode.
  • This is compared against the cavity finesse amplified real PDH error siganl at this stage.
  • Therefore, in frequency noise, the affect of this intensity noise is smaller for higher finesse cavities since cavity finesse only amplifies the PDH signal anf not the scattering noise.

Excess noise estimate

  • I used this measurement taken in 40m with Koji to estimate this noise.
  • This measurement contained a beatnote between IR coupled AUX light and the main laser IR pick-off when X-arm is locked to the main laser and AUX laser is locked to X-arm.
  • So this noise measurement is an upper bound on the total noise in AUX laser frequency when it is locked to the X-arm.
  • I compared this against the noise budget model for AUX PDH loop I have which uses the same control loop as the uPDH box used here.
  • I found a bulge of excess noise below 100 Hz and it seemed to go done as 1/f^2 there. I was reminded by a chat I had with Rana and another professor sometime last year when Rana mentioned scattering noise showing up as "Scatter shelf" looking something like this.
  • So I modeled excess noise as the difference between the noise budget and the measured noise with it extending after 100 Hz with the same roll off as in 10-100 Hz.

Calibration noise budget

  • I took the excess noise measured, converted it to W/rtHz by using current AUX PDH discriminant and photodiode gain, and normalized by the power (9.6mW) to get this noise in RIN/rtHz.
  • Then I assumed that the same RIN would be imparted in the Mariner AUX loop and calculated excess intensity noise at the PDH loop by multiplying the above number with assumed 10mW of incident power to get it in W/rtHz.
  • From here, I fed it to the same input as I feed the shot noise in the loop and calculated the effect in the overall noise budget.
  • For high bandwidth and gain PDH loops required for calibration, this kind of noise would dominate up to a kHz before getting taken over by the residual laser frequency noise.
  • I have again plotted cases for three choices of finesse/mirror transmission. If we used 99.95% reflectivity (1000 ppm transmission, finesse of 3140) we would be fine for most calibration lines except the one around 40 Hz. (assuming drive strength of 1e-13 m everywhere).
  • Otherwise, if possible, we should go for higher finesse. Case (c) plotted here (Page 3), shows that for 99.995% reflectivity (100 ppm transmission, finesse of 31415), we will be fine in all over the range with cavity pole dropping to 63 Hz. This would be really nice of course, if it is possible.
  • So we recommend HR coatings for 1418 nm in the Mariner to be 99.995% reflective (giving total power transmission of 100 ppm).
Attachment 1: AUX_Finesse_Study_With_ECDL.pdf
AUX_Finesse_Study_With_ECDL.pdf AUX_Finesse_Study_With_ECDL.pdf AUX_Finesse_Study_With_ECDL.pdf
  2517   Tue Nov 10 12:46:34 2020 AidanUpdate2um PhotodiodesJPL PD resurrection (cont.)

Looks like the temperature difference between the PD and the shield is relatively small. Even the transients when the heater is applied are order 5K. 

This means that, for quick purposes, the shield RTD is a good proxy for the PD temperature. 

The attached data is the difference between PD and shield RTD from circa 5th-6th February 2020.


Okay - all the steps in the procedure of eLOG 2476 have been verified as working - with the exception of the RTDs in the chamber.

 With regards to taking dark noise spectra at different biases and temperatures, looks like Raymond took spectra with biases of [50, 100, 200, 400, 600, 1000]mV. If no objections, I’ll stick to that number of measurements. 

I’m a bit pushed for time with other stuff. I wonder if the shield RTD is sufficient to run tests on the system? I’ll go back through the data and see how reproducible the relationship between shield temperature and PD temperature is. If it is reliable then in the interests of time, I’m going to forgo re-installing the extra RTDs in the chamber just now.



Attachment 1: temperature_diff_shield_v_PD.png
  2516   Mon Nov 9 15:49:58 2020 AidanUpdate2um PhotodiodesJPL PD resurrection (cont.)

Okay - all the steps in the procedure of eLOG 2476 have been verified as working - with the exception of the RTDs in the chamber.

 With regards to taking dark noise spectra at different biases and temperatures, looks like Raymond took spectra with biases of [50, 100, 200, 400, 600, 1000]mV. If no objections, I’ll stick to that number of measurements. 

I’m a bit pushed for time with other stuff. I wonder if the shield RTD is sufficient to run tests on the system? I’ll go back through the data and see how reproducible the relationship between shield temperature and PD temperature is. If it is reliable then in the interests of time, I’m going to forgo re-installing the extra RTDs in the chamber just now.


  2515   Mon Nov 9 10:08:38 2020 anchalSummaryECDLAUX wavelength finesee requirements in mariner with 1418 nm ECDL (Preliminary)

I have a preliminary calculation to post here. This does not include noise sources from cavity fluctuations and main frequency noise. But it gives some idea about shot noise and frequency noise of AUX laser conttribution to the noise in calibration.

What's included?

  • I have put in measured PZT transfer function of ECDL at ANU upto 25 kHz. Above this point, they did not measure it so I couldn't make it artificially dirty. I just assumed 1/f roll off above (which is definitely incomplete picture).
  • I have included phase effects of cavity FSR in the loop by adding the resonance features as mentioned here.
  • I have added attempted resonance compensation for 1kHz and 2kHz features in the PZT after fitting the data with poles and zeros and iverting them.
  • 10mW of incident AUX light is assumed on the arm cavity.
  • Total 10 mW of combined power at 709 nm is assumed to fall on the beatnote. So AUX light would be frequency doubled for this beatnote.

What's left?

  • Need to add seismic noise and other measured excess noise that come from the cavity motion.
  • Need to add laser frequency noise of main laser, however, it must be small since it is locked to mode cleaner.
  • Need to add digital delay of Red Pitaya or whatever filter would be used for PZT resonance compensation.
  • Need to model PZT transfer function of ECDL above kHz properly. ANU replied that they can't measure it for higher frequencies due to lack of time.
  • Need to do time domain stability analysis. This I haven't been able to do as I have just been using python-controls package as black box to compute impulse and step responses of state space systems. When simply adding easured transfer function data, I couldn't create the state stpace representation for the system. I tried to fit multiple resonances above 2 kHz but couldn't really capture the magnitude of the response well. Maybe I can just assume higher harmonics of the 1kHz and 2kHz resonances?


  • Page 1 is the measured PZT transfer function fo 1900 nm ECDL from ANU along with the modeled 1/f roll off after 25 kHz.
  • Page 2 is the open loop transfer function of the AUX PDH loop for the three different finesse cases studies. Note that the blue curve is hidden beneate the other two curves. Before objections came, I know this is unreal and incomplete, but I have to start somewhere.
  • Page 3 is the calibration noise budget with different colors showing the three finesse cases.
    • This is also incomplete but we can takeaway what the shot noise contribution would look like and initiate a dialogue about the integration time chosen (which is 100s here), the SNR aim chosen (1000 here) and what drive strength would be good enough.
    • From notes of Craig and Gautum, I think we can drive the mirrors at 0.1pm ampltude in the calibration band. From my first and only calibration measurement in 40m, I could drive upto a pm without loosing lock in the cavities.
    • But that was a simple single arm lock and full ifo lock must be more sensitive. So is 0.1 pm drive strength good enough or do we want to aim lower?
Attachment 1: AUX_Finesse_Study_With_ECDL.pdf
AUX_Finesse_Study_With_ECDL.pdf AUX_Finesse_Study_With_ECDL.pdf AUX_Finesse_Study_With_ECDL.pdf
  2514   Fri Nov 6 12:45:21 2020 AidanUpdate2um PhotodiodesPhotodiode testing recovery status

Embellished Chris's PD MEDM screen a bit to illustrate controls in a diagram. The representation of the RELAY SWITCH between the Keithley and the SR560 is a bit off - I think the transimpedance amplifier is switched out as well.

Also - Keithley bright PD sweep output is attached.


Quick update, more detailed update to follow.

  • Laser is working
  • Photodiode sweep with the Keithley shows a sensible dark Current v Voltage plot (when laser is off) - indicating that PD wiring is still intact
  • Laser was aligned onto photodiode (although it took a while aligning to find the signal)
  • Ran a sweep with the Keithley and the laser on - saw another sensible bright Current v Voltage plot (more current than in the dark case)
  • DAQ control still works
  • DAC output is directly providing (unfiltered) 200mV bias

Still to do:

  • Get the SR785 plugged back in
  • Get an SR560 inserted between DAC output and PD bias to low pass filter
  • Investigate why the laser current set point is so noisy
  • Sort out RTD situation inside the chamber
  • Miscellaneous stuff


Attachment 1: MEDM.png
Attachment 2: PD_sweep.png
  2513   Fri Nov 6 08:14:58 2020 AidanUpdate2um PhotodiodesPhotodiode testing recovery status

Quick update, more detailed update to follow.

  • Laser is working
  • Photodiode sweep with the Keithley shows a sensible dark Current v Voltage plot (when laser is off) - indicating that PD wiring is still intact
  • Laser was aligned onto photodiode (although it took a while aligning to find the signal)
  • Ran a sweep with the Keithley and the laser on - saw another sensible bright Current v Voltage plot (more current than in the dark case)
  • DAQ control still works
  • DAC output is directly providing (unfiltered) 200mV bias

Still to do:

  • Get the SR785 plugged back in
  • Get an SR560 inserted between DAC output and PD bias to low pass filter
  • Investigate why the laser current set point is so noisy
  • Sort out RTD situation inside the chamber
  • Miscellaneous stuff
  2512   Thu Nov 5 11:20:45 2020 anchalSummaryECDLTrue PDH Error signal TF and including FSR effects in approximated models

If we use ECDL for auxiliary frequency in 40m and hope to stabilize it up to 1 MHz with digital compensation of PZT, it is important to take into account any phase effect of the nearby FSR at 3.97 MHz. This should ideally be included in the Input Mode Cleaner loop considerations as well. These effects would be more prominent in longer cavities like aLIGO and LISA where FSR is very low and should we attempt to stabilize a laser lock beyond cavity's FSR.

I did a no assumptions calculation for getting a general transfer function fo PDH error signal in units of [W/Hz] assuming 1 W of incident power. This calculation would soon be uploaded here. I'll put down here primary results.

For incident field on a Fabry-Perot cavity (with fsr of \nu_{fsr}), reflected electric field transfer function (unitless) is given by:

\tiny R(\omega) = \frac{-r_1 + (r_2^2 + t_2^2) r_2 e^{-i \omega/\nu_{FSR}}} {1 - r_1 r_2 e^{-i \omega/\nu_{FSR}}}

Then, PDH error signal for a modulation frequency of \Omega at a modulation index of \Gamma, in units of [W/Hz] (i.e. error signal power per Hz of error in laser frequency from cavity resonance) is given by:

\tiny H_{\nu 2P}(\omega) = -\frac{i \pi P_0 J_0(\Gamma)J_1(\Gamma)}{\omega} \Bigl(R(\Omega)R(\omega) - R(0)R(\Omega + \omega) + R^*(\Omega)R(\omega) - R(0)R^*(\Omega - \omega)\Bigr)

after demodulation and low pass filtering. Note this transfer function is a complex quantity as it carries phase information of the transfer function too. The real signal is obtained by multiplying this signal at \omega with e^{i \omega t} and taking the real value of the product.

Having done this, we can see how in the real PDH error signal, there is a low pass at cavity pole, given by - \frac{\nu_{fsr}}{2\pi}log(r_1 r_2) and a notch every fsr. The notch creates a zig-zag in the phase of the tranfer function and has a HWHM same as cavity pole. After this point, I just fitted a ZPK model to the transfer function to obtain a empirically derived model for PDH error signal transfer function. Apart from the cavity pole, this model needs to have resonance and antiresonance features present at each FSR with resonance having a linewidth of cavity pole while anti-resonance having a linewdth of \pi/\nu_{fsr}. Here's how the ZPK model would look like:

 \begin{aligned} z:&\, \pi/\nu_{fsr} \pm n \nu_{fsr} j \\ p:&\, f_p \pm n \nu_{fsr} j, f_p \end{aligned}

I've attached my notebook where I did the fitting analysis and the overlap plot of real PDH error signal TF and the modelled approximation.

Attachment 1: PDHErrorTFModel.pdf
Attachment 2: PDHTFforACavity.ipynb.zip
  2511   Wed Oct 28 14:05:19 2020 ranaSummaryECDLEffects of chosen AUX finesse and source on Calibration requirements

would be easier to achieve there with higher laser powers and higher cavity finesse.

But I haven't attempted that here as we do not know our NPRO PZT's resonance features yet.

I don't know why it would be easier to have higher finesse with longer arms. Something about beam size???

The NPRO PZT TF's are all in the 40m elog - there are many measurements of TF made over the past 10 years. Its like Raiders of the Lost Ark - you have to believe its there while searching.

  2510   Fri Oct 23 12:19:19 2020 anchalSummaryECDLEffects of chosen AUX finesse and source on Calibration requirements

Following up on the last post, here I presented a near back of the envelope calculation of how different choices of AUX cavity finesse and laser source for mariner would affect the prospects of calibration scheme.

Laser sources considered:

As mentioned in the last elog post, here I considered using an NPRO seeded auxiliary laser source (converted to 1418nm by whatever method), ECDL based on ANU design with a modified PDH loop and same ECDl with a digital compensation of PZT resonances. I have taken the residual frequnecy noise of these lasers as the dominant noise source in the calibration scheme. Craig and Gautam in their proposal for SoCal wanted the AUX laser to be locked to the arm cavity in a PDH shot noise limited way. That would be necessary for 4km interferometers and would be easier to achieve there with higher laser powers and higher cavity finesse.

Finesse of 40m Arm Cavity for 1418nm:

Here I considered three cases. First assumes about 3% transmittance of 1418nm in ITM and ETM HR coatings for mariner. This gives a finesse of about 100 and a cavity pole of 18.9 kHz. I believe this is the existing case at 40m. Next we consider transmittance of 0.5% and 0.05% (500 ppm) of 1418nm in ITM and ETM HR coatings for mairner. These cases give finesse of 625 and 6.28k respectively with cavity poles at 3 kHz and 299 Hz respectively.

Page 1: Consideres the case of finesse of 100. The green dashed line shows the amount of drive strength (in m) required at different frequencies if we use ECDL with PZT resonance compensation, to get an SNR of 1000 in 100s of integration time.

Page 2: Same as above but for Finesse of 625.

Page 3: Same as bove but for Finesse of 6280.

Page 4: Comparison of different finesse cases for the ECDL with PZT compensation option. Dashed curves represent requried drive strength (in m) for different cases.

Page 5: Same as above but for NPRO seeded auxiliary laser.

Note: For the NPRO seeded auxiliary laser, we have assumed that the noise of conversion to 1418 nm is similar to noise due to SHG process which is not dominant. There would be an effect of multiplying with a factor ranging form 1-1.5 due to frequency conversion but I have ignored it here for simiplicity. Also, NPRO case is limited in bandwidth due to PZT resonances. We might be able to get away with them using digital compensation like the case study for ECDL. But I haven't attempted that here as we do not know our NPRO PZT's resonance features yet.

Attachment 1: AUX_Finesse_and_Source_Study.pdf
AUX_Finesse_and_Source_Study.pdf AUX_Finesse_and_Source_Study.pdf AUX_Finesse_and_Source_Study.pdf AUX_Finesse_and_Source_Study.pdf AUX_Finesse_and_Source_Study.pdf
  2509   Thu Oct 22 11:19:44 2020 anchalSummaryECDLProspects of using ECDL for Auxiliary laser

We can use Thorlabs SAF1450S2 gain chip to generate 1418 nm light using an ECDL design similar to the one described in Kapasi et al. Optics Express Vol. 28, Issue 3, pp. 3280-3288 (2020) (ANU 2um ECDL design).

PZT Transfer function

I have contacted Disha and Johannes to get the actual measured data for the PZT transfer function of this ECDL design. Fig.5b in their paper plots the transfer function of the PZT. Since, in ECDL PZT directly changes the cavity length, it has a more powerful actuation strength (2 orders of magnitude more) with actuation of 560 MHz.V upto 1 kHz. It however had a very low pole at 1 kHz and two mechanical resonance-antiresonance pairs near 1 kHz and 2 kHz. I modeled a transfer function by eye using Fig.5b of the paper. Page 1 in the attached pdf shows this modelled transfer function.


Next, we need to change the PDH loop for the auxiliary laser lock with the 40m cavity since the PZT has changed. I modelled one from scratch. This simple analog loop's performance is shown in orange in pages 2-5. This loop seemed stable from all the metrics I know, viz: phase margin of about 55 degrees (Page 2), no strong peak in close loop transfer function (page 3), and no remanant oscillations in time domain response (page 4).

I also modeled a similar loop but with digital compensation of the resonance-antiresonance features. This loop is plotted in green on pages 2-5. Both these loops have 300 kHz of bandwidth just by using PZT. I beleive this could be increased but I have not taken into account any saturation of PZT.

ECDL Frequency Noise

From Fig.4. of the paper gives a frequency noise estimate for free running ECDL. They mentioned that a roll-off below 10 Hz was due to their thermal feedback to remain in linear range of their frequency noise emasruement method. I modeled the noise of ECDL hence by
\frac{10^4}{f} + 15\, Hz/\sqrt{Hz}
where the flicker noise contribution is similar to NPRO noise but ECDL has a white noise of 15 Hz/rtHz due to natural linewidth of spontaneous emission or Schawlow-Townes linewidth (with several broadening factors). I think this is an inherent limitation of ECDLs.

Page 5 shows both unsuppressed and suppressed frequency noise estimate for ECDL with the loops mentioned above and current values of NPRO noise are also plotted for comparison.


Attachment 1: AUX_Loop_Study_With_ECDL.pdf
AUX_Loop_Study_With_ECDL.pdf AUX_Loop_Study_With_ECDL.pdf AUX_Loop_Study_With_ECDL.pdf AUX_Loop_Study_With_ECDL.pdf AUX_Loop_Study_With_ECDL.pdf
  2508   Fri Sep 25 14:45:04 2020 StephenDailyProgressCryo vacuum chamberVibration Measurement Equipment and Shield Construction

Came into lab today, with two main goals:

1) Bring Downs equipment for vibration measurements of cryo cooler during next operation

--> DONE, see photo

2) Assemble shields, with two layer scheme using electropolished and mill finish panels.

--> issue: the electropolished shield set was only partially shipped out in Ticket 15101 - unfortunately there was a misinterpretation by Logistics personnel regarding which parts to ship, due to the formatting of the ticket page. I will coordinate shipment of the balance of the panels to the vendor Able Electropolishing for completion of PO S479514

  2507   Wed Sep 23 00:26:59 2020 aaronMiscGenerallab entry

I entered QIL just before Wed Sep 23 00:27:51 2020 to check out and photograph the sprinklers, spent about 20 min looking around the lab and drawing inspiration for in cryo. Wore shoe covers and gloves, touched nothing, sanitized doors after.

  2506   Tue Sep 15 15:16:18 2020 AnchalMiscCleanlinessHEPA Filters turned to maximum

HEPA filters on top of the WOPO table have been turned to High (earlier were at Low).

  2505   Wed Sep 2 08:13:18 2020 StephenDailyProgressCryo vacuum chamberAssembly of QIL Setup and other updates from 2020 Sep 01

2020 Sep 01, StephenA with remote assistance from RaymondR


  • Silicon Mass - Rana had dropped off the Silicon mass in the first room, so I found it when I arrived - thanks!
  • Organization - It was my first time accessing the QIL lab, but everything was pretty well organized and easy to find. All tools for modifications to parts were used in the EE lab which was also well organized. Raymond helped me to figure out where to access things on a few occaisions.
  • Packages - received from Downs Logistics room the Electropolished shield set, a Grainger order with deburring tool and step drill bit, and a McMaster order with a range of bolts - these have all been transported to the lab. Also transported the QIL machined parts that I had received from Machining Solutions to the lab.
    • Koji's Photodiode holders are in the QIL lab ready for pickup.
  • Summary of progress
    • Assembled frame (using torque values from T1100066- #8-32 used 20 in*lb)
    • Assembled brackets to frame
    • Hung silicon mass from music wire
      • Wire is captured under #4-40 SHCS with washers ((using torque values from T1100066- #4-40 used 5 in*lb)
  • Outstanding tasks and questions
    • Did the hang hold?
    • Do we want to have the layered Electropolished and Plain shields during the first installation? Or some sanded state?
    • Assembly requires oversized #8-32 washers which I wasn't able to track down from inventory - these are now on order, along with some more supplies for roughening the surface.

Full Details

  • Assembly work
    • Refer to the DCC - T2000538 - for the videos capturing this assembly effort. I've snagged some screenshots which I've dropped into the attachments.
      • There is a tree catching procedures and other experiment documentation for the QIL Setup at T2000539
  • Issues - there were four issues with the fabricated parts, three of which required small modifications;
    • D2000299-01 small angle rails had threaded holes where there should have been clearance holes for the interface (no issue on big angle rails)
      • modified by drill press to drill out clearance holes at same location
    • D2000308 interface cubes all were threaded only partially through.
      • No action taken, just paid attention and made sure the threads I needed were adequate. Seemed like an offset of only a turn or two, suggesting the CAM program was just a little off (this can happen with tapping, the tap is tapered and the machinist needs to thread deep enough to have the thread major diameter realized through the hole.)
    • D2000307-04 frame upper spacer had threaded holes that were not tapped all the way through.
      • I ran a tap through all of these threads.
    • D2000299-02 large angle rails had threaded holes that didn't pass all of the way through, and we happened to be inserting screws into the wrong side.
      • I ran a screw through these threads, which required a little bit more force than I would have liked, and forcing the screw provided an adequate thread.
        • Note that I anticipate that there will be an issue similar to this, with similar resolution, on the D2000299-01 small angle rails. The shield panels installed on the sides are to be installed from the inside. This can be resolved with a screw coming in from the outside.
    • There was also some inability to access certain screws with the long torque driver, especially if loosening/tightening after putting the frame together.
      • This was managed by use of an L allen key, which of course meant those joints were not torqued to spec. I'm not worried about this compromise.
Attachment 1: photos_cit_qil_lab_cryo_shield_test_assembly_20200901.zip
Attachment 2: T2000538-v1_Part_3_Assembly_of_QIL_Test_Setup_20200901_end_result.jpg
Attachment 3: IMG_7582.JPG
  2504   Fri Aug 14 11:17:04 2020 RaymondUpdateCryo vacuum chamberCooler now operational

Cryopump is up and running. Initial attempts to run the cooler were stymied by an open circuit in the cold head to compressor connection caused by one of the two accessory port fuses (right, circled in attachment 1). The compressor would run but the valve motor wouldn't start in the cooler itself. I extended the spring in the fuse housing (attachment 2) and it seems to have fixed the problem, as now the valve motor starts at the same time the compressor is turned on. Attachment 1 also shows the highly technical cord management procedure done to reduce the trip hazard caused by the compressor power plug. 

User manual recommendations*:

  • Equalization pressure (when not operating): 270-275psi
    • Currently sits at 270psi
  • Operating pressure: 290-330psi
    • Operating at 300psi
  • Insulating vacuum pressure: 1 x 10-3 Torr
    • Vacuum pressure stabilized at 8.5 x 10-1 Torr
      • Edwards diaphragm pump listed ultimate pressure is 1.5 mbar ≈ 1.1 Torr, so either the multirange gauge is malfunctioning or we're getting better backing pressure than expected from the diaphragm pump. A Pirani gauge will be attached to the vacuum space going forward so we'll see how it compares; either way we're above the recommended insulation pressure. The 11" nipple surrounding the coldhead does become cool to the touch during operation, but it does not get cold enough to create condensation.

*Manuals for both the compressor and the cryocooler are linked on the West Bridge wiki manuals page

Attachment 1: IMG-1190.JPG
Attachment 2: IMG-1187.JPG
Attachment 3: IMG-1189.JPG
  2503   Fri Aug 7 11:50:06 2020 RaymondUpdateCryo vacuum chamberTank pumpdown

The diaphragm pump was turned on earlier this week after finally closing up this external adapter tank. Out of an overabundance of caution the tank and cryocooler are supported by the skycrane and a number of posts to prevent it walking off the foam resting pad once the cryocooler is switched on.

All temperature sensors agree with each other within 0.1 K at room temperature yes

Attachment 1: 6A57C6DF-0B58-413E-B9C0-797B14A10CCF_1_105_c.jpeg
  2502   Tue Aug 4 17:08:00 2020 RaymondUpdateCryo vacuum chamber19 pin MIL feedthrough and CTC100 wiring

Used the 19-pin MIL feedthrough to run 4 platinum RTD's and a 25 Ω 100 W resistive heater to the cold head. Attachment 1 is the wiring diagram for the feedthrough and the D-sub connector to the CTC-100 temperature sensor. Attachment 2 shows the three RTDs placed on the cold head. It also shows the thermal anchoring of all lead wires. Attachment 3 shows the RTD attached to the cooler below the cold head using cigarette paper and cryo varnish (stored in the flammables cabinet in QIL).

 The Al block is a premade PT-RTD integrated mounting setup, which was placed on some indium sheet bits and clamped down with a screw and belleville washer. The other two cold head sensors are pressure fit to the cold head by a spring loaded mini dog clamps, and one of the two has some indium underneath the RTD to see if there is any value in doing so going forward with these mounting springs. The glued sensor was attached by painting a thin layer of cryo varnish on the cooler, adding a strip of cig paper, layer of varnish, press in sensor, another strip of paper, paint over all of it with a last thin layer of varnish that reaches beyond the bounds of the paper strips. 

Attachment 1: QIL_TempSensor_MIL19pinWiringDiagram.pdf
Attachment 2: external_tank_RTDs_1.pdf
Attachment 3: F6D69C6A-7168-4D06-B02A-E83CE8AFE524_1_105_c.jpeg
  2501   Fri Jul 24 07:50:00 2020 RaymondUpdateCryo vacuum chamberPrototype shield panels

Picked up the prototype shield panels from Hamilton Metalcraft 7/22 and brought them to QIL. All of the parts are wrapped by part number and in a bin (see attached photo). There are 6 sets of shield panels, but 2 full sets were removed for coating vendors. One full set is as follows (20 parts total):

Part # Quantity
01 2
02 1
03 4
11 2
12 1
13 2
14 2
031 1
032 1
033 1
131 1
132 1
133 1

All component #'s are preceded by 'D2000298-'. 031, 032, and 033 are 03 panels but with hole variations, same goes for 131, 132, and 133 with respect to panel 13  

Attachment 1: MarinerShieldPrototype_parts.jpg
  2500   Fri Jul 24 05:22:30 2020 RaymondLab InfrastructurePD QERound 2 of JPL PD's in lab

Alex dropped off the new round of 2um PD's, they're on the north table accompanied by his data sheet.  

Attachment 1: JPL_PDs_2.jpg
  2499   Fri Jul 24 04:54:57 2020 RaymondLab InfrastructureGeneralMDF repair

First day back (7/15) found the particle board trim w/ powerstrip on the QIL workbench had collapsed. Re-glued and added 4 screws to the middle board where vertical boards from the shelves extended low enough. See attached photos for before, during, and after looks. 

Attachment 1: MDFrepair_before.jpg
Attachment 2: MDFrepair_during.jpg
Attachment 3: MDFrepair_after.jpg
  2498   Fri Jul 24 04:46:51 2020 StephenLab InfrastructureCryo vacuum chamber 

Copper parts picked up July 23rd and brought to QIL, now only waiting on PO# S477874 and the pirani gauge from Koji's bulk JPL order


Torque driver set for QIL setup bolted joints, with range 15 in*oz - 50 in*lb, p/n WIHA 5HYL9, is on order from Grainger, with anticipated delivery in the week of July 20th. Refer to  PO S477925. *update* Tracking Number UPS 1Z19W9330321365493

Cryo connection copper parts PO S475316 will be finished early next week by the machine shop in Torrance, I'll bring them to campus or to Raymond's place (TBD).


  2497   Fri Jul 17 15:54:42 2020 StephenLab InfrastructureCryo vacuum chamber 

Torque driver set for QIL setup bolted joints, with range 15 in*oz - 50 in*lb, p/n WIHA 5HYL9, is on order from Grainger, with anticipated delivery in the week of July 20th. Refer to  PO S477925. *update* Tracking Number UPS 1Z19W9330321365493

Cryo connection copper parts PO S475316 will be finished early next week by the machine shop in Torrance, I'll bring them to campus or to Raymond's place (TBD).

  2496   Tue May 12 10:15:30 2020 aaronUpdatePSOMApreliminary PSOMA layout

I've updated the PSOMA optical layout. I still have some questions on locking, and there are a few additional configurations that we could try. In particular:

  1. Amplifier cavity as a 2-port device (signal and pump mixed before reaching an overcoupled amplifier cavity)
  2. Amplifier cavity as a 4-port device (signal and pump enter the cavity through different ports)
    1. Overcoupled case -- 'signal out' at the same mirror as 'signal in'
    2. Critically coupled case -- 'signal out' through the 'pump in' port

Each of these configurations also has a couple different ways to pickoff an LO for homodyne readout. Shruti and I enumerated these configurations on a zoom whiteboard a couple weeks ago, and I've attached them (the zip contains png).

Chris also mentioned last week that we may run into a frequency-dependent loss in the critically coupled cavity configurations. The pdf I've attached shows a configuration that I think is a minimal modification of the Mach-Zehnder amplifier described in PSOMA. One of the ring cavities is replaced with a tunable steering mirror, and the LO is picked off before the pump reaches the MZ.

In the new diagram, I'm thinking about controlling the following degrees of freedom:

  • Signal phase relative to pump phase
  • Pump frequency relative to amplifier cavity resonance
  • pump intensity
  • MZ relative path length
  • Homodyne mixing angle
  • Signal and pump spatial mode

Some things I'm unsure about:

  • I am not currently controlling signal laser intensity -- can this be done by sending the first order AOM beam to a PD rather than dumping it as shown?
  • Not sure about the right place to place several of the EOM
  • If we have transmission through the cavity's curved mirror, is the associated loss acceptable?

Shruti and I are now tracking our work on git issues in the PSOMA repo.

Attachment 1: whiteboard_configs.zip
Attachment 2: ring_MZ_config.pdf
  2495   Fri May 1 13:27:07 2020 Raymond UpdatePD noiseSb3513 A2P6 2020-02-04 Dark Noise/QE data

Attachment 1 contains the SR785 dark noise measurements at number of PD reverse bias voltages from 77-295K with filenaming convention:

[PD]_drkspec_[date]_[temp]_[input V]_[scan freq]_[FEMTO gain]_[date]_.txt

It also contains the keithley sweeps for QE calculations.

I'm still working out what is wrong with the QE data and how to effectively process the dark noise versus temperature. 

Attachment 1: 20200204_A2P6_77Kto295K.zip
  2494   Thu Apr 16 18:03:22 2020 aaronUpdatePSOMApreliminary PSOMA layout

I put together two PSOMA layouts, one for a bowtie cavity and one for a ring cavity configuration.

I expect there are a number of problems with the layout as I've drawn it, and I note a number of these in the bowtie diagram. Among these

  • Should I put PSOMA updates in the QIL elog, or in Cryo, SUS, or elsewhere? I think QIL has some 1550nm light, as does cryo. Cryo is pretty crowded, although cryo Q has the possibility of moving to the QIL cryostat.
  • What's the best way to get pump and signal -- signal picked off from the pump, or two separate lasers?
  • I expect this layout is rife with errors in how the locking should go. Shruti and I should probably just talk with someone (and each other) about this rather than listing my various uncertainties.
  • Probably also picked off the homodyne LO from a nonideal place
  • I haven't fully considered a number of things like: mode matching, scattered light, LO phase noise and homodyne angle, etc.
Attachment 1: PSOMA_layout.pdf
PSOMA_layout.pdf PSOMA_layout.pdf
  2493   Mon Apr 6 19:01:21 2020 KojiUpdateGeneralWest Bridge flooding Apr 6th

Additional notes:

I did not see anyone in the building.

Attachment 1/2: Our labs have no sticker/paper to indicate any disinfection of the room. (Make sense)

Attachment 3: Most of the basement offices have the notes to indicate disinfection.

Attachment 4/5: Our offices have no notes.

Attachment 1: 20200406151420_IMG_9653.jpg
Attachment 2: 20200406151428_IMG_9654.jpg
Attachment 3: 20200406151459_IMG_9655.jpg
Attachment 4: 20200406151633_IMG_9656.jpg
Attachment 5: 20200406151709_IMG_9658.jpg
  2492   Mon Apr 6 18:38:50 2020 KojiUpdateGeneralWest Bridge flooding Apr 6th

To check the status of all the labs, I went to WB. There was no ongoing water leakage in the labs.

Attachment 1: The subbasement was completely dry.

Attachment 2: Upon the lab inspection, I took PPE from the OMC lab. This was intended to prevent me to pick up anyone's anything and you to pick up my anything.

Attachment 3: The EE shop has no problem

Attachment 4: Cryo Lab. No problem.

Attachment 5: Crackle Lab. No problem, but a lot of dead cockroaches on the floor!

Attachment 6: OMC Lab. No problem.

Attachment 7: C.Ri.Me Lab. Gabriele has already checked the status in the morning. And I found no problem. Didn't bother to turn on the light.

Attachment 8: CTN Lab. No problem.

Attachment 9: QIL Lab. The floor was mostly dry. Did someone wipe the floor?

Attachment 10: Some water drip was found in front of the workbench.

Attachment 11: It comes from the ceiling.

Attachment 12: Left a trash box to catch future possible leak.

Attachment 13/14: TCS Lab. No problem found.

Attachment 15: As per Aidan's request, the instruments were moved to the North-East area of the room to avoid future possible leak.

Attachment 1: 20200406143251_IMG_9618.jpg
Attachment 2: 20200406143856_IMG_9621.jpg
Attachment 3: 20200406143932_IMG_9622.jpg
Attachment 4: 20200406144014_IMG_9623.jpg
Attachment 5: 20200406144119_IMG_9626.jpg
Attachment 6: 20200406143837_IMG_9620.jpg
Attachment 7: 20200406144413_IMG_9633.jpg
Attachment 8: 20200406144522_IMG_9635.jpg
Attachment 9: 20200406144648_IMG_9639.jpg
Attachment 10: 20200406144730_IMG_9643.jpg
Attachment 11: 20200406144752_IMG_9644.jpg
Attachment 12: 20200406144942_IMG_9645.jpg
Attachment 13: 20200406145125_IMG_9646.jpg
Attachment 14: 20200406145127_IMG_9647.jpg
Attachment 15: 20200406145347_IMG_9652.jpg
  2491   Mon Apr 6 18:35:48 2020 KojiSummaryGeneralWest Bridge flooding Apr 6th

West Bridge flooding Apr 6th due to rain in the night

Looks like the first responder was Calum. The attached photos were sent from him.


Attachment 1: image2.jpeg
Attachment 2: image1.jpeg
  2490   Mon Mar 9 13:13:02 2020 KojiLab InfrastructureHVACHVAC work concluded for today

Last Friday, I found the HEPA units on the squeezer table were not on. I turned them on at "SLOW".


  2489   Fri Feb 28 13:36:32 2020 Ian MacMillanSummaryLab MonitoringItem lending: Particle Counter from OMC Lab to QIL

Still trying to figure out how to set up the particle counter remotely. The current particle count is 576.

Particle counts over time
Feb. 28 at 12:30pm 576
Feb. 28 at 5:00pm 594
Mar. 2 at 8:30am 393
Mar. 2 at 11:30am 650

Note: the particle count is the number of particles detected over 0.3um size.

  2488   Thu Feb 27 14:26:52 2020 KojiSummaryLab MonitoringItem lending: Particle Counter from OMC Lab to QIL

Item lending as per Ian's request: Particle Counter from OMC Lab to QIL

The current particle class of the room was measured to be 800.

The particle counter went back to the OMC lab on Aug 10, 2020.

Attachment 1: P_20200227_134755_vHDR_On.jpg
  2487   Tue Feb 25 15:12:22 2020 Raymond, Chris, Koji, Chub, AidanLab InfrastructureHVACHVAC work concluded for today

Facilities workers replumbed the water lines feeding the air handler units in the QIL. Tomorrow they plan to come back for about an hour to insulate the lines. We'll keep the tables wrapped up until that's complete.

  2486   Mon Feb 24 18:41:06 2020 RaymondLab InfrastructureHVACTable Shielding for 2020-02-25 HVAC repairs

[Raymond, Chris, Koji, Chub, Aidan, Duo]

Both tables were surrounded with plastic shielding in preparation for the HVAC maintenance/repairs to be performed tomorrow, Tuesday 24/02/2020. The IR Lab cryostat roughing and turbo pumps were turned off to avoid overheating while inside their new plastic cocoon. Sticky floor mats were placed in the oil-slick area in front of the flammables cabinet to improve safety for all foot traffic in the area, as well as to mitigate the urge to boogie while working in the lab. 

Attachment 1: QIL_HVACshielding1_2020-02-24.JPG
Attachment 2: QIL_HVACshielding2_2020-02-24.JPG
Attachment 3: QIL_HVACshielding3_2020-02-24.JPG
Attachment 4: QIL_HVACshielding4_2020-02-24.JPG
  2485   Thu Feb 13 11:04:47 2020 Ian MacMillanLab InfrastructureHVACRecent temperature fluctuations

Here is the data for the last week. The temperature for the Northside is noticeably higher than the Southside. This is probably a calibration error in the North sensor because there is not a noticeable temperature difference when walking across the room. I would guess the south sensor is more accurate in its overall temperature reading because it feels more like 22 degrees than 24.

Attachment 1: HVAC-B265B.pdf
Attachment 2: Jan_2020_Qil_Temp_Measurments.zip
  2484   Tue Feb 4 14:18:24 2020 Ian MacMillanLab InfrastructureHVACRecent temperature fluctuations

Got data from the HOBO temperature recorders. The temperature has been fluctuating between about 80 degrees F to 60 degrees F. The Temperature has been fluctuating wildly in the last month that the recorder was on. I will restart the HOBO monitors to get more current data.

*Note the HOBO app changed how it reports the time data. The updated code is attached which uses the AM and PM 12 hour .csv file generated by the app.

Attachment 1: HVAC-B265B_20200402.pdf
Attachment 2: Qil_Lab_Temperature.zip
Attachment 3: HVAC-B265B.pdf
  2483   Fri Dec 20 22:26:19 2019 KojiUpdatePD QEPD TEC driver / A2P6 aligned / Lens moved

The QEs were measured at 293K, 239K, 232K, and 293K again. The cooling was provided by the PD TEC.  At each temperature, the incident power was changed from 30uW to 1mW to see the dependence of the QE on the incident power to check the possible saturation.

The QE was 79~81% (the window T=96.6% was already compensated). I'm not 100% sure this 1% variation in the plateau is real or due to insufficient calibration of the REF PD.
The REF PD was calibrated at 1mW at 100mA injection current to the laser.

No obvious saturation was observed.

We can cool the PD with LN2 and we should make a careful alignment of the beam at each temperature.

Attachment 1: Sb3513_A2P6_DarkCurrent_293K.pdf
Attachment 2: Sb3513_A2P6_DarkCurrent_239K.pdf
Attachment 3: Sb3513_A2P6_DarkCurrent_232K.pdf
Attachment 4: Sb3513_A2P6_DarkCurrent_293K_2.pdf
Attachment 5: Sb3513_A2P6_DarkCurrent_Comparison.pdf
Attachment 6: 191220_3513A2P6.zip
  2482   Fri Dec 20 21:58:14 2019 KojiUpdatePD QEPD TEC driver / A2P6 aligned / Lens moved

== Currently, A2P6 is aligned ==

1) I've brought another TEC driver fro the PD temp control. This unit was borrowed from the 2um ECDL setup. Eventually, we need to return this to ECDL. (Attachment 1)
The PID loop of the TEC control works. But it is not well optimized yet. If you change the target temp too quickly, the TEC out seemed oscillating. Watch the TEC out carefully and change the temp setpoint slowly.
So far I have tried to cool the thermister up to 30kOhm (~232K) and I_TEC was 0.33A.
I did not try further. I felt it was better to cool the PD base for further trial.

2) A part of the alignment study, the beam is aligned to A2P6. Also, the lens position was investigated, and I decided to move the lens ~1 inch away from the window.  (Attachment 2)
    In fact, this allowed us to insert the power meter between the lens and the window. 

Attachment 1: P_20191220_192440_vHDR_On.jpg
Attachment 2: P_20191220_180929_vHDR_On.jpg
  2481   Fri Dec 20 13:20:53 2019 AidanSummaryPD QEQE results from A2P6 (500um) and A2P2 (1mm)

The QE measurements from the first couple of photodiodes are attached below.

  • plot_JPL_diode_results.m - A2P6 analysis
  • plot_JPL_A2P2_diode_results.m - A2P2 analysis

QE = [I_photocurrent]/[P_PD] * h *nu/e

P_PD = Power incident on photodetector = 0.966*power_incident on cryo window

Power incident on cryo window = F(voltage on reference PD)

Attachment 1: PC_DC_v_T.pdf
Attachment 2: A2P2_001_test.pdf
Attachment 3: PC_DC_v_T.pdf
Attachment 4: A2P6_001_test.pdf
Attachment 5: plot_JPL_diode_results.m
% load JPL data
f0 = dir('*dark*.txt');
f1 = dir('*photo*.txt');
f2 = dir('*cond*.txt');

% get temperature vs time
tempList = [];
pList = [];
for ii = 1:numel(f2)-1
... 102 more lines ...
Attachment 6: plot_JPL_A2P2_diode_results.m
close all 
clear all
% load JPL data
f0 = dir('*dark');
f1 = dir('*bright*');

% get temperature vs time
tempList = [];
refPDList = [];
for ii = 1:numel(f1)
... 113 more lines ...
  2480   Mon Dec 16 18:16:31 2019 RaymondDailyProgressPD QECryostat wiring fixes

Opened the cryostat to resolder the heater and re-wrap the thermal anchor for the sample RTD and PD connection. All connections are working as expected at room temperature. A2 is still in the sample mount. 

  2479   Fri Dec 13 01:51:15 2019 RaymondDailyProgressPD QECryostat wiring fixes
  • Monitored an autorun of the photocurrent for A2P6 from about 77K-300K and adjusted the laser alignment as necessary. I've not yet looked through the data, as I was keen on fixing the assorted wiring issues in time to pump down for tomorrow. Chris found the laser power issue rooted in my sub-optimal diode-collimator mating connection, ie I didn't fully plug back in the laser. I've learned my lesson and will simply shield the fiber with Al foil for future nitrogen pouring endeavours .
  • Opened the cryostat and confirmed the pin assignments (see attached list).
  • Cathode 2 was shorting on the thermal anchors for the PD plug. I re-wrapped the quad twist wire around the bobbins and this has fixed the issue, though I did not see any portion of the quad twist wires that was missing the Formvar insulation. 
  • The DT-670 silicon diode thermometer was busted, this was likely the case from the start and should have been checked (by me) prior to wiring. I've replaced this diode with another 4-lead platinum resistor, it is plugged into input 3 on the CTC100 (still labeled 'PDdiode'). 
  • The thermistor on A2P6 is still intact, showing about 1.9 kΩ at room temp

Still to do:

  • Move the radiation shield PT RTD to the outer shield
  • Re-wire the baseplate pins in the DSUB9 connection to the CTC100 temp controller (switched V+ and V-)
  • Check feasibility of adding 25Ω heater to the sample mount to stabilize the temp of the PD rather than measuring while sweeping 


Attachment 1: 18_32malepinsinairAssigned.pdf
  2478   Thu Dec 12 02:10:37 2019 Raymond, ChrisDailyProgressPD QEPD test dry run

(Raymond, Chris)

  • Baseplate and shield RTD temperature sensors are working, and digitally recorded.  The temperature sensing diode on the PD mount seems to be unconnected.
  • We made dark current vs bias measurements on the P6 and P2 elements at room temperature.  We also checked P3 but it appears to be unconnected.  Then, after realigning the P6 element, we measured its photocurrent vs bias at room temperature.  The photocurrent and dark current data looks much like what Koji previously reported at room temperature.
  • Going from 77 K to room temperature, the beam had to be moved downward to correct the alignment.  The magnitude was about 1 clockwise turn of the vertical alignment knob on the steering mirror.  The horizontal shift was very small.
  • Armed with this knowledge, we re-wetted the cryostat with LN2, bringing the baseplate and shield temperatures down into the 80 K range.  Our plan was to let it warm up overnight while running an automated test series of dark current and photocurrent vs bias and temperature.  We offset the vertical alignment by a half turn, in hopes of having the PD well aligned near the expected QE sweet spot of 150-200 K.
  • The procedure for adding LN2 to the cryostat involves temporarily disconnecting the fiber coupler.  Sadly, it appears in our haste we did not fully recover the alignment after this step, or else the calibration somehow shifted by a lot.  The photocurrent is at the 0.1 mA level instead of the ~1 mA we expect.  Accordingly, this run will primarily be useful as a test of the automation, as an indication of dark current vs temperature, and as a very rough, qualitative indication of QE vs temperature.
  • The script is running on qil-ws1 in a screen session.  It can be interrupted with screen -RAad autorun to connect to the session, followed by Ctrl-C to kill the script.
  • In the morning, we need to decide whether to open the cryostat and fix the wiring, or to repeat this run with tighter control over the alignment.
  2477   Wed Dec 11 19:50:25 2019 ranaHowtoCryo vacuum chamberProcedure to record photodiode output vs temperature

don't we also want to record the dark noise spectrum as a function of T and V_Bias ? I would guess that the dark noise doesn't always scale with dark current at low frequencies since its probably more like a random walk than shot noise.

  2476   Wed Dec 11 15:53:48 2019 Aidan, Chris, RaymondHowtoCryo vacuum chamberProcedure to record photodiode output vs temperature

[Aidan, Chris, Ray]

  1. Add LN2 to cryo-chamber
  2. Turn on heater to 25W
  3. Wait for LN2 to boil off
  4. Turn on the REFPD bias switch, and enable a 200 mV bias on the JPL PD (set C4:TST-FM13_OFFSET to 200)
  5. Switch the readout to the transimpedance amplifier (relay control output from C4:TST-FM15 at 30k ct)
  6. Turn on the laser
    1. Power up the LDC201C and set it for half-power output (50 mA), with the modulation output from C4:TST-FM12 at zero
    2. Set C4:TST-FM12_OFFSET to 1 to go up to full power
  7. Monitor the PD output on oscilloscope and adjust horizontal and vertical alignment of laser beam so that the output of the PD is maximized (also adjust TIA gain if needed so the output is not saturated)
  8. Switch the readout to the Keithley (relay control output from C4:TST-FM15 at zero)
  9. Record bright PD response
    1. Turn on laser (set C4:TST-FM12_OFFSET to 1)
    2. Scan bias voltage and record the PD response using the source meter (scripts are located in $HOME/JPL_PD/scripts on qil-ws1)
      1. ./runsweep.py triggers the Keithley to sweep the bias (range of sweep is defined in the script)
      2. ./getdata.py FILENAME downloads data from the Keithley and writes it to FILENAME
    3. Record REF PD reading
    4. Record RTD resistance
  10. Record dark PD response
    1. Turn off laser (set C4:TST-FM12_OFFSET to -1)
    2. Repeat steps 9-2 through 9-4
  11. Switch the readout to the transimpedance amplifier (relay control output from C4:TST-FM15 at 30k ct)
  12. Measure dark noise vs bias (using either the SR785 or the cymac, TBD)
  13. Wait two minutes and repeat steps 6 through 12 (there is a script ./autorun.sh which continuously repeats steps 9 and 10)
  14. Continue until RTD reaches room temperature (approximately 60-90 minutes).
ELOG V3.1.3-